OpenStudy (anonymous):

solve the quadratic equation x^2 -11x-12=0

OpenStudy (anonymous):

@hbaldon ?

OpenStudy (anonymous):

x^2 = 11 x+12

OpenStudy (anonymous):

(x-12) (x+1) = 0

OpenStudy (anonymous):

(x-11/2)^2-169/4 = 0

OpenStudy (anonymous):

find the ans

OpenStudy (anonymous):

x = -1

OpenStudy (anonymous):

again, factor out x^2-11x-12=0 and solve for x. :D

OpenStudy (anonymous):

it has the same procedure from your 1st question.

OpenStudy (anonymous):